OpenStudy (anonymous):

Solve for x: 3|x - 3| + 2 = 14 No solutions x = -1, x = 8.3 x = 0, x = 7 x = -1, x = 7

OpenStudy (anonymous):

No Solution

OpenStudy (accessdenied):

3|x - 3| + 2 = 14 <--- we subtract from both sides, 2 3|x - 3| = 12 <--- divide off the 3 |x - 3| = 4 <--- two possible cases. x - 3 = +4 and x - 3 = -4 generally there is no solution when we have, say, an absolute value equal to a negative number. we should never have |1| = -1 or something of the sort.

OpenStudy (anonymous):

So , Its NO SOLUTION RIGHT

OpenStudy (accessdenied):

|x - 3| = a positive 4 o.o the situation i am referring to is like this: |x - 3| = negative 4 imagine the graph of the absolute value if that helps. it has a vertex on the x-axis and is alwayys positive. you cannot intersect something always positive with a number y = -4, but certainly y=+4 runs through it.

OpenStudy (anonymous):

its a yes or no question ;-; im very confused non of that even is a answer

OpenStudy (accessdenied):

we are meant to give an explanation for the answer, not give them blindly. ;x this is where i left off: x - 3 = +4 and x - 3 = -4 if you understood the work before it and continue through, you should get a final answer by solving each equation.

OpenStudy (anonymous):

Ikr @Courtneydyrdek98